a particle move with velocity v1 for time t1 and v2 for time t2 along a straight line .The magnitude of average acceleration is..
 

Grade:11

2 Answers

Abhinav
8 years ago
use
av acceleration= change in velocity/total time

change in velocity=v2-v1
total time= t1 +t2
soni kumari
3 years ago
avgerage acc. = delta V | T
                    
                    CHANGE IN VELOCITY |TOTAL TIME
 
      =  V- V1  |T+T2  ans. 
v stands for VELOCITY     t stands for TIME